I don't get this.
How would I solve it.
Please help!
The salary of an employee was divided as follows: 1/5 for electric bill, 1/2 for food, and 1/6 for transportation. How much of the salary did he save?

Answer by rfer(16322)

1/5+1/2+1/6=
find common denominator to add
6/30+15/30+5/30=26/30=13/15 spent
2/15 saved